OpenStudy (anonymous):

Which of the following expressions is not equal to 81x^3y^5 ?

OpenStudy (anonymous):

choices??

OpenStudy (anonymous):

yea im trying lol

OpenStudy (anonymous):

sowy

OpenStudy (anonymous):

A.3^2 xy^2 * 3^2 x^2 y^3 B.3xy(27xy)xy3 C.3^3 x^3 *3^2 y^5 D.3^4 * x^5 * y^2 * x^-2 y^3

OpenStudy (anonymous):

D is the answer

OpenStudy (anonymous):

3*3*3*3=81 x^5*x^-2=x^3 y^2*y^3=y^5